# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ 3 Operator systems and realizations
The memory system of a finitely correlated state $\omega$ is by definition a finite-dimensional operator system (cf. Appendix A), which means that it can be represented as a finite-dimensional subspace of some algebra $\mathcal{B}(\mathcal{H})$. However, the underlying Hilbert space $\mathcal{H}$ does not need to be finite-dimensional, and indeed there are examples of operator systems that can only be embedded into infinite-dimensional $C^{*}$-algebras (see [34, Thm. 2.4] together with the fact that the group $C^{*}$-algebra of a free group is infinite-dimensional). Moreover, an arbitrary finitely correlated state is not necessarily $C^{*}$-finitely correlated (in fact, already in the abelian case [11] there are counterexamples, but we don’t know of properly non-commutative examples). Now, it is true that $C^{*}$-finitely correlated states are weak-*-dense in all translation invariant states [8, Prop. 2.6]. Thus, we cannot distinguish between $C^{*}$-finitely correlated states and a generic translation invariant state by measuring only finitely many observables up to precision $\varepsilon>0$. But since the memory size needed to realize these approximating $C^{*}$-finitely correlated states will usually diverge with the size of the region where the observables are supported, we are also interested in extending the algorithm directly to the case of a general finitely correlated state.
@@
The order we consider in the following is constructed by taking $\Psi_{n}(X)$ to be positive if $X$ is positive. This construction can also be dualized which gives us a minimal and a maximal order. This will implicitly be used in the proof of Lemma 3.1 and is commented on in Remark 3.3. The order gives rise to an order norm $\|\Psi_{X}\|_{\Psi_{\mathds{1}}}$ (see Appendix A) that in the case of a $C^{*}$ order yields back the $C^{*}$ norm. In addition to this generalized operator norm, we will also need an analogue of the the Hilbert-Schmidt norm for the estimates in Section 4.3. Since we cannot guarantee the memory systems to be represented on a finite-dimensional Hilbert space we cannot directly use the Hilbert-Schmidt norm. However, we can use the maps $\Psi_{X}$ defined in Equation (13) to define an inner product. For $X\in\mathcal{A}_{[1,t^{*}]}$ set $\xi_{X}(i)=\operatorname{Tr}\left(XX_{i}\right)$ and $X^{\prime}=\sum_{i}\Psi_{X}(X_{i})X_{i}$. Then clearly $\Psi_{X}(Y)=\operatorname{Tr}\left(X^{\prime}Y\right)$ and $\xi_{X^{\prime}}(i)=\Psi_{X}(X_{i})$, so $\xi_{X^{\prime}}=\ddot{\Omega}\xi_{X}$ where $\ddot{\Omega}_{ij}=\omega(X_{i}\otimes X_{j})$ (‘Ömega’) is $\Omega$ written in coordinates. Furthermore, $\Psi_{X}=\Psi_{Z}$ if and only if $\xi_{X^{\prime}}=\xi_{Z^{\prime}}$ if and only if $X^{\prime}=Z^{\prime}$. We then set
@@
showing that $\Psi_{2,r\mathds{1}+S(X)}$ is not a positive functional. By definition of the order norm this shows $r\leq\|\Psi_{X}\|_{\Psi_{\mathds{1}}}$ and thus the desired norm inequality (cf. also Remark 3.3). ∎
Remark 3.2.
Replacing $\mathcal{C}_{R}$ by $\mathbb{M}_{n}(\mathcal{C}_{R})\cong\mathbb{M}_{n}\otimes\mathcal{C}_{R}$ (similarly for $\mathcal{C}_{L}$) and $\Psi$ by its $n$-th amplification $\Psi_{n}$, we can apply the above proof mutatis mutandis. This yields the inequality
$\|\cdot\|_{n,\Omega}\leq\sqrt{n}\|\cdot\|_{{\Psi_{n}}(\mathds{1})}$
for the norm on $\mathbb{M}_{n}(\mathcal{C}_{R})$.
The inequalities in Lemma 3.1 hold for the norms defined on the memory system of the abstract (minimal) realization. Consider the restriction of $U^{\intercal}$ to $V$, $\Psi_{X}\mapsto U^{\intercal}\Psi_{X}$, which is invertible. It induces an operator system structure on $\mathbb{C}^{m}$ that is completely order isomorphic to the order structure of the regular realization and it holds that $U^{\intercal}\Psi_{\mathds{1}}=e$ as defined in Section 2.2.1. Also note that $\lVert(\mathrm{id}_{\mathbb{M}_{n}}\otimes U^{\intercal})\Psi_{n,X}\rVert_{2}=\lVert\Psi_{X}\rVert_{n,\Omega}$.
@@
In fact, the matrix order on the memory system $V$ is not unique. The construction in [8] mentioned in the beginning of this section can be applied to both sides of the chain. Dualizing the order on one side then gives a matrix order on the other side and the two need not coincide. However, every positive element in the directly constructed one is also positive in the one obtained through dualizing and in fact every admissible matrix order making the generating map $\mathbb{E}$ completely positive has to lie between the two orders. We may thus call them minimal and maximal and since the cones of the minimal matrix order are smaller than those of the maximal matrix order, the corresponding order norms satisfy the reverse inequality. This has also been observed by I. Todorov [35] (and possibly others). The proof of Lemma 3.1 works for the minimal matrix order for the leftmost inequality and implicitly uses the maximal matrix order for the rightmost. Therefore, the bounds are valid for the order norm of any admissible matrix order.
Remark 3.4.
@@
Every bipartite state between two $C^{*}$-algebras gives rise to operator system structures analogous to Remark 3.3. Assuming that this operator system is finite-dimensional, it is also sufficient to consider a finite-dimensional subspace of the algebra, which can be assumed to be an operator (sub) system. If this operator subsytem can be represented on a finite-dimensional Hilbert space, we can again use the Hilbert-Schmidt inner product to define a norm on the space of ‘correlation functionals’ that satisfies the bounds as in Lemma 3.1. This might be useful to obtain bounds in other settings. We are not aware of this being mentioned in the literature explicitly in this context. However, the results of [36, 35, 37] are very close to this.
Remark 3.5.
@@
We also note that a quotient realization of a $C^{*}$-finitely correlated state can be given the structure of an abstract operator system, which is not necessarily the minimal or the maximal order, but is inherited by the usual matrix order on $\mathcal{B}$. $\mathcal{V}=\mathcal{W}/{\mathcal{W}\cap\tilde{\mathcal{W}}^{\perp}}$ can also be given the structure of an operator system, with unit $e^{\prime}=L\mathds{1}_{\mathcal{B}}$ and cones $C_{n}:=L(\mathbb{M}_{n}(\mathcal{W})^{+})$.
@@
By “push-forward” via the invertible map $M$, the observable realization also gets an operator system structure and $M$ automatically becomes unital and completely positive. The map $R$ identifying $\mathcal{V}$ with $\mathcal{V}^{\prime}\subset\mathcal{W}$, see Lemma 2.9, then is a (unital) complete order isomorphism. $\mathcal{V}$ then inherits the Hilbert-Schmidt norm from $\mathcal{B}$, which we denote as $\lVert\cdot\rVert_{2}$ without ambiguity. Then, for $x\in\mathbb{M}_{n}(\mathcal{V})$ we have $\lVert x\rVert_{(Le^{\prime})^{(n)}}\leq\lVert Rx\rVert_{\mathds{1}_{\mathbb{M}_{n}(\mathcal{B})}}\leq\lVert x\rVert_{2}$ by contractivity of $L$ and the standard inequality between Schatten norms.
We thus have
@@
$\lVert\mathrm{id}_{\mathbb{M}_{n}}\otimes U^{\intercal}\mathcal{F}R\rVert_{\mathds{1}_{\mathbb{M}_{n}}\otimes Le^{\prime}\rightarrow 2}\leq\lVert\mathrm{id}_{\mathbb{M}_{n}}\otimes\mathrm{id}_{\mathbb{C}^{m}}\rVert_{e\rightarrow 2}\lVert\mathrm{id}_{\mathbb{M}_{n}}\otimes U^{\intercal}\mathcal{F}R\rVert_{\mathds{1}_{\mathbb{M}_{n}}\otimes Le^{\prime}\rightarrow{\mathds{1}_{\mathbb{M}_{n}}\otimes e}}\leq\sqrt{n}.$ (52)
4 State reconstruction for FCS
@@
Consider the corresponding map $\Omega_{(\cdot)}:(A,A^{\prime})\mapsto\Omega_{A}(A^{\prime})$ as defined in Equation (16) and its observable realization $(\mathbb{C}^{m},e,\mathbb{K},\rho)$ as in Proposition 2.5. We can reconstruct an approximation of the realization parameters from an estimate $\hat{\Omega}_{(\cdot)}$ obtained from estimating the expectation values of $\omega_{t}$. Here, we need to impose that the estimate $\hat{\Omega}$ has at least rank $m$, and we take its $m$-truncated singular value decomposition, meaning that we only retain the first $m$ singular values and singular vectors. This means that $\hat{U}$ is the matrix of the first $m$ left singular vectors of $\hat{\Omega}$.
Definition 4.1 (Spectral state reconstruction).
@@
4.1 Empirical estimates of realization parameters
@@
Therefore using Lemma B.2 and the inequality $\|A\|_{2\rightarrow 2}=\lVert A\rVert_{\infty}\leq\|A\|_{2}$ and the fact that $\sigma_{m}(\hat{\Omega})=\sigma_{m}(\hat{U}^{\intercal}\hat{\Omega})$,
@@
Moreover, given bases $\{X_{\ell}\}_{\ell=1}^{m_{1}}$, $\{A_{j}\}_{j=1}^{d^{2}}$, $\{Y_{i}\}_{i=1}^{m_{2}}$ of $\mathcal{C}_{R}$, $\mathcal{A}$ and $\mathcal{C}_{L}$, respectively,
$\|\Omega_{(\cdot)}\|_{2}$ $\leq\sqrt{\sum_{i=1}^{m_{2}}\sum_{j=1}^{d^{2}}\sum_{l=1}^{m_{1}}|\omega(Y_{i}\otimes A_{j}\otimes X_{\ell})|^{2}}\leq\sqrt{\operatorname{Tr}[\omega_{2t^{*}+1}^{2}]}\leq 1\,.$ (68)
@@
Then using $\|\tau\Omega\|_{2}=\sqrt{\sum_{i=1}^{m_{1}}{\omega}(X_{i})^{2}}\leq 1$ and Lemma B.2,
@@
Then,  (59), (62), (63) directly follow from perturbation bounds for singular values, which are collected in Appendix B, and in particular Lemma B.5. For these bounds we closely follow the treatment of previous works in the classical setting [12, 18, 19]. ∎
4.2 Error propagation bound
With $(\mathbb{C}^{m},\mathcal{K},e,\rho)$ being the observable realization Proposition 2.5, in the following we will assume that $\mathcal{V}$ is an operator system with order unit $e^{\prime}$ and $M:\mathcal{V}\rightarrow\mathbb{C}^{m}$ is an invertible linear map such that $Me^{\prime}=e$ and that $\rho^{\prime}=\rho M$ is a positive functional. It holds that $\rho^{\prime}(e^{\prime})=1$ and that the map ${M}^{-1}{\mathbb{K}}{M}$ is unital, in the sense that ${M}^{-1}{\mathbb{K}}_{\mathds{1}_{\mathcal{A}}}{M}e^{\prime}=e^{\prime}$. We further assume that it is completely positive. Of course, we can always take $\mathcal{V}$ to be the vector space of functionals $\Psi_{A}$, $A\in\mathcal{A}_{R}$, whose positive elements at level $n$ are $\Psi_{n}(A)$, $A\in\mathbb{M}_{n}(A)^{+}$. The following error parameters will be used in the analysis: given $\widetilde{e}$, $\widetilde{\rho}$, $\widetilde{\mathbb{K}}_{A}$ and $\hat{M}$ as in Equations (39), (40), (41) and (42). respectively, where $\hat{U}$ and is chosen as in Definition 4.1,

### v3

@@ B.2 Operator systems and finitely correlated states
The memory system of a finitely correlated state $\omega$ is by definition a finite-dimensional operator system (cf. Section B.1), which means that it can be represented as a finite-dimensional subspace of some algebra $\mathcal{B}(\mathcal{H})$. However, the underlying Hilbert space $\mathcal{H}$ does not need to be finite-dimensional, and indeed there are examples of operator systems that can only be embedded into infinite-dimensional $C^{*}$-algebras (see [46, Thm. 2.4] together with the fact that the group $C^{*}$-algebra of a free group is infinite-dimensional). Moreover, an arbitrary finitely correlated state is not necessarily $C^{*}$-finitely correlated (in fact, already in the abelian case [17] there are counterexamples, but we don’t know of properly non-commutative examples). Now, it is true that $C^{*}$-finitely correlated states are weak-*-dense in all translation invariant states [14, Prop. 2.6]. Thus, we cannot distinguish between $C^{*}$-finitely correlated states and a generic translation invariant state by measuring only finitely many observables up to precision $\epsilon_{HS}>0$. But since the memory size needed to realize these approximating $C^{*}$-finitely correlated states will usually diverge with the size of the region where the observables are supported, we are also interested in extending the algorithm directly to the case of a general finitely correlated state.
Let $\Psi_{n}\coloneqq\mathrm{id}_{\mathbb{M}_{n}}\otimes\Psi$, and $V_{n}\coloneqq\{\Psi_{n}(X),X\in\mathbb{M}_{n}(\mathcal{A}_{R})\}$. The order we consider in the following is given by the cones
$C_{n}\coloneqq\overline{\{\Psi_{n}(X)|X\in\mathbb{M}_{n}(\mathcal{A}_{R}),X\geq 0\}},$
@@
where the closure is in weak-* topology, with units $e_{n}=\Psi_{n}(\mathds{1})$. It is immediate to check that this construction gives a bona-fide operator system. The construction can also be dualized which gives us a minimal and a maximal order. This will implicitly be used in the proof of Lemma B.6 and is commented on in Remark B.8. The order gives rise to an order norm $\|\Psi_{X}\|_{\Psi_{\mathds{1}}}$ (see Appendix B.1) that in the case of a $C^{*}$ order yields back the $C^{*}$ norm. In addition to this generalized operator norm, we will also need an analogue of the the Hilbert-Schmidt norm for the estimates in Section C.3. Since we cannot guarantee the memory systems to be represented on a finite-dimensional Hilbert space we cannot directly use the Hilbert-Schmidt norm. However, we can use the maps $\Psi_{X}$ defined in Equation (42) to define an inner product. For $X\in\mathcal{A}_{[1,t^{*}]}$ set $\xi_{X}(i)=\operatorname{Tr}\left(XX_{i}\right)$ and $X^{\prime}=\sum_{i}\Psi_{X}(X_{i})X_{i}$. Then clearly $\Psi_{X}(Y)=\operatorname{Tr}\left(X^{\prime}Y\right)$ and $\xi_{X^{\prime}}(i)=\Psi_{X}(X_{i})$, so $\xi_{X^{\prime}}=\ddot{\Omega}\xi_{X}$ where $\ddot{\Omega}_{ij}=\omega(X_{i}\otimes X_{j})$ (‘Ömega’) is $\Omega$ written in coordinates. Furthermore, $\Psi_{X}=\Psi_{Z}$ if and only if $\xi_{X^{\prime}}=\xi_{Z^{\prime}}$ if and only if $X^{\prime}=Z^{\prime}$. We then set
@@
showing that $\Psi_{2,r\mathds{1}+S(X)}$ is not a positive functional. By definition of the order norm this shows $r\leq\|\Psi_{X}\|_{\Psi_{\mathds{1}}}$ and thus the desired norm inequality (cf. also Remark B.8). ∎
Remark B.7.
Replacing $\mathcal{C}_{R}$ by $\mathbb{M}_{n}(\mathcal{C}_{R})\cong\mathbb{M}_{n}\otimes\mathcal{C}_{R}$ (similarly for $\mathcal{C}_{L}$) and $\Psi$ by its $n$-th amplification $\Psi_{n}$, we can apply the above proof mutatis mutandis. This yields the inequalities
$\sigma_{m}(\Omega)\|\cdot\|_{{\Psi_{n}}(\mathds{1})}\leq\|\cdot\|_{n,\Omega}\leq\sqrt{n}\|\cdot\|_{{\Psi_{n}}(\mathds{1})}$
for the norms on $\mathbb{M}_{n}(V)$.
The inequalities in Lemma B.6 hold for the norms defined on the memory system of the abstract (minimal) realization. Consider the restriction of $U^{\intercal}$ to $V$, $\Psi_{X}\mapsto U^{\intercal}\Psi_{X}$, which is invertible. It induces an operator system structure on $\mathbb{C}^{m}$ that is completely order isomorphic to the order structure of the regular realization and it holds that $U^{\intercal}\Psi_{\mathds{1}}=e$ as defined in Section A.2.1. Also note that $\lVert(\mathrm{id}_{\mathbb{M}_{n}}\otimes U^{\intercal})\Psi_{n,X}\rVert_{2}=\lVert\Psi_{X}\rVert_{n,\Omega}$.
@@
In fact, the matrix order on the memory system $V$ is not unique. The construction in [14] mentioned in the beginning of this section can be applied to both sides of the chain. Dualizing the order on one side then gives a matrix order on the other side and the two need not coincide. However, every positive element in the directly constructed one is also positive in the one obtained through dualizing and in fact every admissible matrix order making the generating map $\mathbb{E}$ completely positive has to lie between the two orders. We may thus call them minimal and maximal and since the cones of the minimal matrix order are smaller than those of the maximal matrix order, the corresponding order norms satisfy the reverse inequality. This has also been observed by I. Todorov [47] (and possibly others). The proof of Lemma B.6 works for the minimal matrix order for the leftmost inequality and implicitly uses the maximal matrix order for the rightmost. Therefore, the bounds are valid for the order norm of any admissible matrix order.
Remark B.9.
@@
Every bipartite state between two $C^{*}$-algebras gives rise to operator system structures analogous to Remark B.8. Assuming that this operator system is finite-dimensional, it is also sufficient to consider a finite-dimensional subspace of the algebra, which can be assumed to be an operator (sub) system. If this operator subsytem can be represented on a finite-dimensional Hilbert space, we can again use the Hilbert-Schmidt inner product to define a norm on the space of ‘correlation functionals’ that satisfies the bounds as in Lemma B.6. This might be useful to obtain bounds in other settings. We are not aware of this being mentioned in the literature explicitly in this context. However, the results of [48, 47, 49] are very close to this.
Remark B.10.
@@
We also note that a quotient realization of a $C^{*}$-finitely correlated state can be given the structure of an abstract operator system, which is not necessarily isomorphic to the the minimal or the maximal order, but is inherited from the usual matrix order on $\mathcal{B}$. $\mathcal{V}=\mathcal{W}/({\mathcal{W}\cap\tilde{\mathcal{W}}^{\perp}})$ can also be given the structure of an operator system, with unit $e^{\prime}=L\mathds{1}_{\mathcal{B}}$ and cones $C_{n}:=L(\mathbb{M}_{n}(\mathcal{W})^{+})$. By “push-forward” via the invertible map $M$, the observable realization also gets an operator system structure and $M$ automatically becomes unital and completely positive. This operator system, given by the cones $MC_{n}$ is admissible in the sense above. Indeed, the minimal operator system in the observable realization is given by the cones
$C^{\min}_{n}\coloneqq\overline{\{\mathbb{K}^{n}_{X}e|X\in\mathbb{M}_{n}(\mathcal{A}_{[1,n]}),X\geq 0,n\geq 1\}},$
and the units $e_{n}$. The maximal operator system is still given by the units $e_{n}$, but the cones are
$C^{\max}_{n}\coloneqq\{\rho\mathbb{K}^{n}_{Y}|Y\in\mathbb{M}_{n}(\mathcal{A}_{[-n,0]}),Y\geq 0,n\geq 0\}^{*}.$
Using the notation of Proposition A.4, we have $\mathbb{K}^{t}_{X}e=M\mathbb{K}^{\prime t}_{X}e^{\prime}=L\mathcal{E}^{t}_{X}(\mathds{1}_{\mathcal{B}})$, and since $\mathcal{E}^{t}_{X}(\mathds{1}_{\mathcal{B}})\geq 0$ if $X\geq 0$, $\mathbb{K}^{t}_{X}e\in C^{\min}_{n}$ implies $\mathbb{K}^{t}_{X}e\in MC_{n}$. On the other hand, if $Y\geq 0$ and $\mathcal{E}^{t_{1}}_{X}(\mathds{1}_{\mathcal{B}})\geq 0$ (so that $\mathbb{K}^{t_{2}}_{X}e\in MC_{n}$),
$\rho\mathbb{K}^{t_{1}}_{Y}\mathbb{K}^{t_{2}}_{X}e=\rho^{\prime}\mathbb{K}^{\prime t_{1}}_{Y}\mathbb{K}^{\prime t_{2}}_{X}e^{\prime}=\rho_{0}\mathcal{E}^{t_{1}}_{Y}|_{W}\mathcal{E}^{t_{2}}_{X}(\mathds{1}_{\mathcal{B}})\geq 0,$ (83)
and thus $\mathbb{K}^{t_{2}}_{X}e\in C_{n}^{\max}$.
@@
The map $R$ identifying $\mathcal{V}$ with $\mathcal{V}^{\prime}\subset\mathcal{W}$, see Lemma A.11, then is a (unital) complete order isomorphism. $\mathcal{V}$ inherits the Hilbert-Schmidt norm from $\mathcal{B}$, which we denote as $\lVert\cdot\rVert_{2}$ without ambiguity. For $x\in\mathbb{M}_{n}(\mathcal{V})$ we have $\lVert x\rVert_{(e^{\prime})^{(n)}}\leq\lVert Rx\rVert_{\mathds{1}_{\mathbb{M}_{n}(\mathcal{B})}}\leq\lVert x\rVert_{2}$ by contractivity of $L$ and the standard inequality between Schatten norms. We also have that $U^{\intercal}\mathcal{F}Re^{\prime}=e$, and for the norms induced by the cones $C_{n}$,
@@
$\lVert\mathrm{id}_{\mathbb{M}_{n}}\otimes U^{\intercal}\mathcal{F}R\rVert_{\mathds{1}_{\mathbb{M}_{n}}\otimes e^{\prime}\rightarrow 2}\leq\lVert\mathrm{id}_{\mathbb{M}_{n}}\otimes\mathrm{id}_{\mathbb{C}^{m}}\rVert_{\mathds{1}_{\mathbb{M}_{n}}\otimes e\rightarrow 2}\lVert\mathrm{id}_{\mathbb{M}_{n}}\otimes U^{\intercal}\mathcal{F}R\rVert_{\mathds{1}_{\mathbb{M}_{n}}\otimes e^{\prime}\rightarrow{\mathds{1}_{\mathbb{M}_{n}}\otimes e}}\leq\sqrt{n},$ (84)
where we used that the (82) (in particular, the second inequality) holds for the norms of any admissible matrix order on the memory system and, thus $\lVert\mathrm{id}_{\mathbb{M}_{n}}\otimes\mathrm{id}_{\mathbb{C}^{m}}\rVert_{\mathds{1}_{\mathbb{M}_{n}}\otimes e\rightarrow 2}\leq\sqrt{n}$, and that $U^{\intercal}\mathcal{F}R$ is a unital complete order isomorphism.
Appendix C State reconstruction for FCS
@@
Consider the corresponding map $\Omega_{(\cdot)}:(A,A^{\prime})\mapsto\Omega_{A}(A^{\prime})$ as defined in Equation (45) and its observable realization $(\mathbb{C}^{m},e,\mathbb{K},\rho)$ as in Proposition A.5. We can reconstruct an approximation of the realization parameters from estimates $\hat{\Omega},\hat{\Omega}_{(\cdot)},\widehat{\Omega(\mathds{1})},\widehat{\tau{\Omega}}$ obtained from estimating the expectation values of $\omega_{t}$. Here, we need to impose that the estimate $\hat{\Omega}$ has at least rank $m$, and we take its $m$-truncated singular value decomposition, meaning that we only retain the first $m$ singular values and singular vectors. This means that $\hat{U}$ is the matrix of the first $m$ left singular vectors of $\hat{\Omega}$.
Definition C.1 (Spectral state reconstruction).
@@
C.1 Empirical estimates of realization parameters
@@
Therefore using Lemma G.2 and the inequality $\|A\|_{2\rightarrow 2}=\lVert A\rVert_{\infty}\leq\|A\|_{2}$ and the fact that $\sigma_{m}(\hat{\Omega})=\sigma_{m}(\hat{U}^{\intercal}\hat{\Omega})$,
@@
Moreover, given bases $\{X_{\ell}\}_{\ell=1}^{d_{\mathcal{A}}^{2s}}$, $\{A_{j}\}_{j=1}^{d_{\mathcal{A}}^{2}}$, $\{Y_{i}\}_{i=1}^{d_{\mathcal{A}}^{2s}}$ of $\mathcal{C}_{R}$, $\mathcal{A}$ and $\mathcal{C}_{L}$, respectively,
$\|\Omega_{(\cdot)}\|_{2}$ $\leq\sqrt{\sum_{i=1}^{d_{\mathcal{A}}^{2s}}\sum_{j=1}^{d_{\mathcal{A}}^{2}}\sum_{l=1}^{d_{\mathcal{A}}^{2s}}|\omega(Y_{i}\otimes A_{j}\otimes X_{\ell})|^{2}}\leq\sqrt{\operatorname{Tr}[\omega_{2t^{*}+1}^{2}]}\leq 1\,.$ (100)
@@
Then using $\|\tau\Omega\|_{2}=\sqrt{\sum_{i=1}^{d_{\mathcal{A}}^{2s}}{\omega}(X_{i})^{2}}\leq 1$ and Lemma G.2,
@@
Then,  (91), (94), (95) directly follow from perturbation bounds for singular values, which are collected in Appendix G, and in particular Lemma G.5. For these bounds we closely follow the treatment of previous works in the classical setting [20, 27, 28]. ∎
C.2 Error propagation bound
With $(\mathbb{C}^{m},\mathcal{K},e,\rho)$ being the observable realization from Proposition A.5, in the following we will assume that $\mathcal{V}$ is an operator system with order unit $e^{\prime}$ and $M:\mathcal{V}\rightarrow\mathbb{C}^{m}$ is an invertible linear map such that $Me^{\prime}=e$ and that $\rho^{\prime}=\rho M$ is a positive functional. It holds that $\rho^{\prime}(e^{\prime})=1$ and that the map ${M}^{-1}{\mathbb{K}}{M}$ is unital, in the sense that ${M}^{-1}{\mathbb{K}}_{\mathds{1}_{\mathcal{A}}}{M}e^{\prime}=e^{\prime}$. We further assume that it is completely positive. Of course, we can always take $\mathcal{V}$ to be the vector space of functionals $\Psi_{A}$, $A\in\mathcal{A}_{R}$, whose positive elements at level $n$ are $\Psi_{n}(A)$, $A\in\mathbb{M}_{n}(A)^{+}$. The following error parameters will be used in the analysis: given $\widetilde{e}$, $\widetilde{\rho}$, $\widetilde{\mathbb{K}}_{A}$ and $\hat{M}$ as in Equations (70), (71), (72) and (73). respectively, where $\hat{U}$ and is chosen as in Definition C.1,